### v1

# RL in the Wild: Characterizing RLVR Training in LLM deployment

Abstract
Large Language Models (LLMs) are now widely used across many domains. With their rapid development, Reinforcement Learning with Verifiable Rewards (RLVR) has surged in recent months to enhance their reasoning and understanding abilities. However, its complex data flows, and diverse tasks pose substantial challenges to RL training systems, and there is limited understanding of RLVR from a system perspective. To thoroughly understand the system challenges introduced by RLVR, we present a characterization study of RLVR tasks in our LLM deployment. Specifically, we investigate the distribution and variation trends of workloads across different RL tasks across training steps. We identify issues such as GPU idling caused by skewed sequence length distribution, inefficient parallel strategies in dynamically varying workloads, inefficient data management mechanisms, and load imbalance. We describe our observations and call for further investigation into the remaining open challenges. Furthermore, we propose PolyTrace benchmark suite to conduct evaluation with realistic workloads, a practical use case validates that PolyTrace benchmark suite exhibits 94.7% accuracy.
1 Introduction
The rapid advancement of LLMs (e.g., DeepSeek[35], OpenAI o1[46], Gemini[65]) has fundamentally transformed the landscape of academia and industry. This exponential growth in model complexity has been accompanied by increasingly sophisticated training paradigms, evolving from traditional Supervised Fine Tuning (SFT) to more nuanced approaches that better improve model capabilities. The computational demands of training these massive models have necessitated the development of distributed training systems[61, 55, 83, 85] capable of handling huge scales of data and computation.
Within this broader context of LLM development, RLVR has emerged as a critical methodology for post-training optimization. Unlike conventional pretraining[9] and SFT[64, 4, 47] phases, RLVR introduces a fundamentally different computational paradigm where models interact with environments, receive feedback signals, and iteratively refine their behavior through policy optimization algorithms such as Proximal Policy Optimization (PPO) [57] and Group Relative Policy Optimization (GRPO) [58]. This approach has proven essential for enhancing their reasoning capabilities[62], leading to widespread adoption in both research and industry.
To enhance the performance of RLVR training systems, A thorough workload analysis is essential for comprehending challenges and uncovering opportunities. However, existing workload analysis conducted for other LLM tasks, such as pretraining[26, 34] and serving[76, 52, 73], do not apply to RLVR training systems. This is primarily due to the divergent characteristics and requirements of RLVR:
Workload Pretraining Serving RL
Prior Work [23, 11, 17] [76, 73, 74] -
[…]
#Model 1 $1\sim$4
#Stage 2 2 8+
Static Data
Heterogeneity
[…]
Table 1: Comparison of Pretraining, Serving, and RL workload characteristics. RL workloads are the most complex, involving multiple stages, multiple models, dynamic data generation, and a tightly coupled software stack.
(1) Longer step span. In RLVR, a single training step often takes on the order of minutes and, in some cases, can exceed one hour. By contrast, other LLM training tasks typically complete a step in several seconds, while an inference step (prefill and decode) is usually under one second. This significantly increases fault recovery costs, hinders real-time monitoring.
(2) Complicated workflow. In RLVR, a single training run requires multiple models to collaborate according to the algorithmic logic, and is often divided into several stages, such as rollout, inference, and training. By contrast, other LLM training tasks typically involve only one model and a single training stage[47, 53]. This introduces challenges to resource allocation and handling failure propagation across stages.
(3) Unknown sequence length. In prior LLM training tasks, the workload was known in advance, enabling offline reordering and scheduling policy selection[11]. In RLVR, however, the workload—much like that of LLM serving—cannot be predetermined and exhibits dynamics and abrupt shifts. This leads to suboptimal parallelization strategies and difficulty in applying appropriate workload-aware scheduling.
(4) Complex software stack. Previously, LLM tasks could be executed within a single training or inference framework, such as Megatron[61], FSDP[83], vLLM[28], or SGLang[86]. In RLVR, however, the training framework must be built atop existing training and inference frameworks, and additionally incorporate software for environment interaction, such as sandboxes[3] and MCP-based[39] programs.
To bridge this gap, we present PolyTrace and an in-depth analysis of RLVR training systems based on our operational experiences across multiple production tasks. To the best of our knowledge, this represents the first comprehensive study of RLVR workloads in real-world deployment scenarios. Our analysis draws upon extensive traces collected from large-scale RLVR training jobs, encompassing detailed profiling data, resource utilization metrics, and system performance logs. Our further key findings can be summarized as follows:
  • Various and long-tail sequence length distributions. Different RL tasks exhibit substantial variations in input and output sequence length distributions. In terms of output length, mathematics and image understanding tasks generate extremely long output sequences, with average lengths reaching 9,839.3 and 5,911.6 tokens, respectively, whereas video understanding and searching tasks exhibit average output lengths of only 166.5 and 48.7 tokens. The pattern for input lengths is markedly different: due to the presence of video information, video understanding tasks have an average input length of 3,123.1 tokens, while mathematics tasks require merely 126.5 tokens. Meanwhile, long-tail distributions are prevalent across RL tasks. For instance, in tool-use tasks, the 90th percentile versus the maximum input length differs by up to 18×, while the 90th percentile versus the maximum output length reaches approximately 5×.
  • Dynamically varying output length during training. On one hand, the variation trends in output length are model-dependent: larger models typically do not exhibit pronounced output length changes, whereas smaller models demonstrate significant sequence length variation trends. On the other hand, different tasks also show markedly distinct variation patterns. For example, searching tasks exhibit decreasing output lengths as training progresses, while mathematics tasks show increasingly longer outputs.
  • Fluctuating system performance during training process. In large-scale RL training tasks, performance varies dramatically across training steps. For instance, in image understanding tasks, the best performance can reach 341 tokens per GPU per second (TGS), while the worst performance drops to merely 0.8 TGS. This instability stems from diverse sequence lengths and inappropriate parallelization strategies.
Based on our characterization study, we identify several challenges and insights in RL training systems. Firstly, workload-aware request scheduling strategies are essential. The diverse sequence length distribution characteristics significantly impact RL training system performance, and simple scheduling strategies such as First Come First Serve (FCFS) cannot achieve satisfactory performance in RL training. SLO-aware scheduling strategies(e.g., prefill and decoding disaggregation[89]) are not profitable in RL training. secondly, inefficient data transmission and management implementations. To improve framework usability, RL training frameworks often adopt single-controller designs, but in certain tasks, this design creates single-point bottlenecks that lead to degraded training performance or even failure. thirdly, severe load imbalance. The unbalanced sequence length distributions in RL training result in uneven, uncertain, and dynamic sequence length during the training stage; current static parallelization strategies and coarse-grained load balancing algorithms exhibit poor performance. Other challenges, such as expensive scheduling overhead in inference frameworks, coarse-grained and inefficient GPU memory management, and uncertain tool invocation latencies, indicate that RL training systems require further design and optimization. The benchmark suite serves as a crucial tool for uncovering system insights in new cases and enabling researchers to validate their optimization methods effectively. To this end, we propose PolyTrace and PolyTrace benchmark suite to uncover system insights and comprehensive evaluations with realistic workloads. A practical use case validates that PolyTrace benchmark suite exhibits 94.7% accuracy.
We believe the observations and insights from our work are broadly applicable. Based on our analysis, we highlight the profound impact of workload distribution on the performance of RL training systems. Consequently, our traces and benchmark suite are publicly available at https://github.com/Zhou-jiecheng/PolyTrace to facilitate further research and development in this area.
2 Background
2.1 LLM Reinforcement Learning
Figure 1: Overview of the RL Training Workflow.
RL Training Workflow. As illustrated in Figure 1, unlike pre-training and SFT, RLVR typically comprises three stages and multiple LLMs. The workflow can be divided as follows:
1. Rollout. The actor model, analogous to an agent model in the RL setting, performs autoregressive generation[70] starting from an initial state. Each generated token constitutes an action that transitions to a new state. Once generation terminates, the resulting sequence is treated as a trajectory for subsequent stages. This stage is usually executed by inference engines (e.g., vLLM[28], SGLang[86], and LMdeploy[82]). The actor model may also invoke external tools (e.g., image processing[88], web search[24]) at this stage.
2. Inference. This stage may involve multiple models. At a minimum, it includes forward passes of the actor model and a reference model, typically orchestrated by the training framework[61, 83, 55]. To obtain the reward for the actor model, different RL tasks have diverse choices, commonly falling into three categories: (a) Rule-based rewards. Directly match the response produced by the actor model against predefined rules to assign a reward score. This is suitable for tasks with readily verifiable correctness, such as coding or mathematics. (b) Reward model-based rewards: When the reward cannot be formalized, a trained reward model—often a fine-tuned base LLM—provides the reward score via a forward pass. (c) Judge-model-based rewards: To avoid the substantial cost of training a reward model, some methods directly use a stronger LLM (often a larger one) to evaluate the response and produce a reward score. Unlike the previous cases, this step is commonly executed via an inference engine. The critic in the GRPO algorithm is a relative value calculated based on a set of responses from the same prompt, whereas in the PPO algorithm, it is provided through a critic model.
3. Training. The actor model and, in PPO, the critic model are trained to update model parameters. This stage resembles pretraining and supervised fine-tuning (SFT) in that it is handled directly by LLM training frameworks such as Megatron[61] or FSDP[83]. Training LLMs efficiently at scale necessitates various system innovations, such as state sharding optimizers[54], meticulous model placement using data[31], pipeline[42], and tensor parallelisms.
Because training objectives differ across tasks [62], models "react" differently to each task, resulting in highly diverse workloads. In such a multi-stage pipeline, workload characteristics exert a complex and pronounced impact on the system.
RL training system The RL training systems can be primarily categorized into synchronous and asynchronous. The core distinction lies in whether there exists staleness in the model version in the rollout stage of the RL training process.
1. Synchronous RL training. As illustrated in Figure 2 (a), in Synchronous RL training, each phase strictly respects data dependencies; consequently, rollout and training are executed with identical parameter versions. While straightforward, this strategy incurs severe GPU under-utilization when the workload is highly skewed, limiting its scalability.
2. Asynchronous RL training: Recent algorithmic studies[10] relax these dependencies by allowing the rollout phase to lag behind training by up to a maximum permitted staleness. This pipelining reduces GPU idle time (Figure 2 (b)). Moreover, the two phases exhibit distinct compute characteristics: rollout is usually memory-bound due to long generation lengths, whereas training is compute-bound. Exploiting this disparity enables the use of heterogeneous hardware[90], improving training efficiency and lowering cost.
Figure 2: Comparison of Synchronous and Asynchronous RL Training. We omit additional stages (e.g., inference) to better highlight the fundamental distinction between synchronous and asynchronous RL training.
2.2 PolyTrace Overview
Motivated by prior discussions and the evident gap in publicly available traces, we introduce PolyTrace to fill this void. As shown in Table 4, PolyTrace comprises workloads from seven representative reinforcement learning (RL) tasks, including four open-source RL training tasks and three large-scale RL training tasks used in model development. Each item entry captures the following fields:
  1. 1.
    Step. The global training step to which the record belongs.
  2. 2.
    Input length. To preserve anonymity for certain internal datasets, we disclose only the input length for each input.
  3. 3.
    Output length. The length of the model output associated with each prompt.
  4. 4.
    Type. The task category of the workload record.
Unlike serving traces [51, 74], RL training workloads intentionally omit arrival timestamps. In RL training, samples within the same step can be treated as arriving concurrently. PolyTrace serves two complementary purposes: on one hand, our comprehensive trace is designed to catalyze systems research on RL training optimization, on the other hand, it facilitates rigorous, apples-to-apples evaluation of RL frameworks under diverse tasks.
3 Characterize RL Training in the Wild
In this section, we present a thorough analysis of RL training workloads collected from RL Training logs during model production. Our study includes three kinds of RL tasks: image understanding for model A in Section 3.1, mathematics and tool use (Section 3.3) for model B (Section 3.2). The training configurations are detailed in Table 2
3.1 Vision-Task Workloads
Figure 3: CDF of Image Understanding Task input and output length distribution
Figure 4: Scatter of Image Understanding Task input and output length joint distribution
Figure 5: Heatmap of Image Understanding Task prompt and outputs, we randomly sample 64 prompts for readability
Figure 6: The workloads distribution across training steps of Image Understanding Task.
Figure 7: Similarity of distribution across taining steps of Image Understanding Task. The metric in the figure is 1 minus the Jensen–Shannon (JS) divergence.
Figure 8: Figure (a) shows the average sequence lengths of correct and incorrect answers. Figure (b) depicts the performance at each mini training step during the training stage, where the performance metric is Tokens per GPU per Second (TGS). In RL training, once the rollout completes the generation of data for a global batch, the algorithm partitions this global batch into several mini-batches (8 in this case).
Model Arch Param GPUs Placement Train Par Rollout Par
A MoE 235B 256 Colocate
FSDP, SP=4
EP=16
B MoE 235B 128 Colocate
TP=4, ETP=1
EP=8, PP=8
TP=8
Table 2: Training Configuration. Both of them use FP8 parameters during the rollout stage. We did not scale the training to a very large size due to the poor scalability of RL training (Section 4.4).
[…]
Joint input–output structure. We further examine the input–output joint distribution. As shown in Figure 4, the distribution exhibits a banded pattern, corresponding to the input-distribution pattern in Figure 3. There is no pronounced correlation between input and output lengths, only around an input length of roughly 2k do long-tail outputs become less likely. The potential factor is that input length is largely decoupled from prompt difficulty and will not affect the model’s reasoning length. Across steps, the joint distributions remain strongly similar, reflecting the difficulty distribution of prompts sampled from the dataset is similar across steps.
Prompt-correlated output lengths. In RL training, algorithms typically perform multiple stochastic samples (rollouts) for the same prompt, producing a set of outputs. As illustrated in Figure 5, each prompt is sampled 16 times; darker colors denote shorter outputs, while blank cells indicate that the algorithm filtered out low-quality training samples. We observe strong similarity in output lengths across samples for the same prompt. This implies that per-sample output length is not fully random: for the same prompt, the model tends to produce outputs of comparable length, likely because the token budget required to solve the same underlying problem is similar across samples. As a result, we recommend scheduling at the prompt granularity rather than per sample. Also, it is possible to first run a preliminary sampling pass to obtain sample output lengths, then schedule other samples accordingly, rather than relying on an additional model to predict the output length[90].
Similar temporal distribution. In Figure 6, we further examine how the input and output length distributions evolve across training steps. As training progresses, the medians of input and output lengths exhibit only minor jitter without clear temporal trends. However, output-length distributions can fluctuate more substantially across steps: for example, at step 13, the 95th percentile is around 16k, whereas at steps 41 and 12 it reaches roughly 25k. Moreover, Figure 7 shows that output-length distributions across steps are strongly similar, while input distributions lack clear cross-step similarity. This is primarily due to the carefully designed prompt sampling strategy and the diversity of the datasets. Last but not least, RL often operates with very few epochs—indeed, for Model A, the epoch count can be less than one—rendering epoch-based optimizations (e.g., cross-epoch data reuse) less effective and of limited generality. In Figure 8 (a), we can observe that samples with correct answers generally have longer sequence lengths than those with incorrect answers, across training steps. which may indicate tha,t as training progresses, more correct responses will lead to increased sequence lengths.
Training performance volatility. As illustrated in Figure 8 (b), during the training stage, performance exhibits significant fluctuations, with the maximum TGS reaching 341 while the minimum TGS is only 0.8 ( 400x difference). Based on our in-depth analysis, there are two primary contributing factors: (1) The input-output length distribution (Figure 3) and sample filtering strategy[37] result in substantial variations in sequence lengths across mini training steps. For instance, at the step where TGS is 0.8, the training samples have a sequence length of only 435, whereas at the step where TGS reaches 341, the training samples have a sequence length of 31,280. Longer sequence lengths entail greater computational workload, thereby overlapping the overhead introduced by communication. (2) Static parallelization strategies fail to achieve satisfactory performance. Throughout this training process, we consistently employed the same parallelization configuration with sequence parallel = 4, which severely impacts training throughput. When sequence lengths are short, using sequence parallel[20] introduces additional communication overhead instead. This observation underscores the significance of an online sequence-length-aware parallelization strategy in RL training.
3.2 Mathematics-Task Workloads
Figure 9: CDF of Mathematics Task input and output length distribution
Figure 10: Scatter of Mathematics Task input and output length joint distribution
Figure 11: The workloads distribution across training steps of Mathematics Task.
To strive for a deeper understanding of the characteristics of different workloads in RL training during model production, we further conduct comprehensive analysis of Model B.
[…]
Stability of distributions over training. We next examine how these distributions evolve during training in Figure 11. The input and output length distributions for math show no pronounced temporal trend; however, we observe a very gradual increase in output length. This likely reflects improving reasoning ability as RL progresses, which elongates the chain of thought. Due to the very large parameter scale and the strong capabilities imparted by pretraining and SFT, large models under RL do not exhibit the pronounced lengthening seen in smaller models as displayed in Section 3.4.
3.3 Tool-use Task Workloads
Figure 12: CDF of Tool-use Task input and output length distribution
Figure 13: Scatter of Tool-use Task input and output length joint distribution
Figure 14: The workloads distribution across training steps of Tool-use Task.
Figure 15: The workloads distribution across training steps of Tool-use Task.
Beyond the mathematics task, we further conduct a comprehensive analysis of the tool-use task in Model B.
[…]
Clustered joint input–output distributions. In Figure 13, we analyze the joint distribution of input and output lengths for the tool-use task. Unlike the largely uncorrelated patterns observed in the previous two tasks, the joint distribution here clusters by the number of dialogue turns, forming roughly five groups, with a few isolated outliers around the cluster peripheries. This corresponds to the concentration of turns between 2~6 and most samples share similar length profiles in Figure 12. However, the presence of isolated outliers complicates RL training system optimizations.
Turn-related input/output length distributions. We further study how length distributions relate to dialogue turns and training steps. In Figure 15 (a), input length scales almost linearly with the number of turns, because each additional turn prepends prior turns’ inputs and outputs to the LLM context, and per-turn reasoning chains and tool-return lengths are themselves relatively concentrated (see Figure 12 (b) and Figure 12 (d)). In Figure 15 (b). Interestingly, output length also depends on turn count: most long-tail outputs occur at a single turn, suggesting that, absent tool calls, the model may produce longer chains of thought. Once tool calls are involved, output length drops markedly, with the median decreasing from about 639 tokens to roughly 130 tokens, indicating that the model can quickly integrate context and make a tool-calling decision with far fewer tokens.
Step-related input length dynamics. We next analyze how length distributions evolve with training steps in the tool-use task. Unlike the previous two tasks, the input length exhibits a distinct regime shift. As shown in Figure 14 (a), there is a sudden drop around step 30, after which the series transitions from stable to sustained fluctuations. The reason shown in Figure 14 (c), as training progresses, tool calls become more efficient and less frequent, reducing the number of dialogue turns and thus the total input length. Figure 14 (b) shows that output lengths remain nearly unchanged throughout training, implying a stable deliberation process when the model decides whether and how to invoke tools.
3.4 Other-task Workloads
Figure 16: Cumulative distribution function (CDF) plots of the input and output length for each task.
Figure 17: Violin plots of the input and output length for each task.
Figure 18: The trend of workload variation among different tasks.
To further analyze the characteristics of workload distribution in RL training, we collect and analyze the workloads of four RL tasks with relatively smaller models.
Various sequence length distribution among RL tasks. As shown in Figure 16, within a single iteration, the distributions of different RL training tasks exhibit distinct characteristics. In terms of output length, programming tasks produce very long outputs, with an average of 9,329, whereas the video understanding and searching tasks average only 166.5 and 48.7, respectively. The pattern is reversed for input lengths: due to the presence of video information, the video understanding task has an average input length of 3,123.1, while the other tasks do not exceed 1k. Looking at the standard deviation of outputs, the coding task is the most uneven, with a standard deviation reaching 8,651.5, followed by the mathematics task at 595. By contrast, the multimodal task shows the most uneven input distribution, mainly because video length varies. As shown in Figures 16 and 17, both the mathematics and programming tasks display pronounced heavy-tailed behavior in their output distributions, whereas the other two tasks do not. In contrast, the searching task exhibits a pronounced heavy tail in its input-length distribution, which is not evident for the other three tasks. This is primarily because searching tasks require multiple external tool calls to retrieve information and passing it to the model, the number of tool invocations varies, leading to long and highly uneven inputs. We emphasize the need to comprehensively consider these workloads when optimizing RL training systems.
Markedly different sequence length variation patterns. Compared to LLM RL training, the input and output length evolve continuously in relatively smaller models, with both the trend and magnitude of change differing across task types as shown in Figure 18. For example, in mathematics and programming tasks, output length tends to increase over time. In agent tasks, output length gradually decreases, whereas in the video understanding task, the input and output lengths tend to stabilize. This phenomenon arises because different RL tasks prioritize distinct model capabilities. For instance, coding and mathematics tasks emphasize strengthening reasoning, which encourages the model to produce longer chains of thought; accordingly, output lengths tend to increase as training progresses. In contrast, the searching task focuses on decision-making proficiency. As training advances, the model becomes more adept at tool use, thereby shortening the chain of thought required to complete a problem. Compared with large-scale models in Section 3, small models exhibit significantly greater workload variability. We highlight the necessity of workload awareness when performing dynamic parallelism strategy selection in RL system optimization.
The widespread use of mixed datasets to boost overall capability further increases distributional diversity[67, 13], adding to the challenges of RL framework design and optimization.
4 Fine-grained RL Training System Analysis
This section analyzes system implications of RL training tasks as introduced in Section A (the first four tasks). Firstly, we provide a time breakdown analysis of the different components of RL tasks in Section 4.2, including fine-grained and coarse-grained. Moreover, we analyze the GPU memory usage in Section 4.3. Lastly, we conduct a scalability study (Section 4.4) and hyper hyperparameter study (Section 4.5) across RL tasks.
4.1 Experiment Setup
Algorithm: The detailed configurations of the four open source algorithms we analyze are presented in Section A.
Hardware configuration: Our experiments are conducted on two clusters. Cluster A comprises 16 machines (128 GPUs), where each machine is equipped with 8 NVIDIA A800 80GB GPUs interconnected via 400GB/s NVLink, with an inter-machine bandwidth of 200Gbps. Cluster B consists of 2 machines, each equipped with 8 NVIDIA H800 80GB GPUs interconnected via 400GB/s NVLink, with an inter-node bandwidth of 1600Gbps. Our experiments utilize the following software versions: CUDA 12.4[44], PyTorch 2.6.0, NCCL 2.21.5[45], Verl 0.3.1, Sglang 0.4.6 and vLLM 0.8.5.
4.2 Time Breakdown
Figure 19: Coarse-grained time breakdown across different RL tasks.
Figure 20: Per-rank time breakdown of video understanding task in training stage.
The workloads of different tasks vary significantly according to PolyTrace, which directly affects the time breakdown. Analyzing these divergent breakdowns can provide guidance for optimizing bottlenecks in the RL training system.
Coarse-grained time breakdown. Stage-wise time shares in RL training are task-dependent. The rollout stage typically takes the lion’s share, and the breakdown fluctuates to varying extents. As shown in Figure 19, we observe substantial cross-task variation in the percentage of total time attributable to each stage. In most tasks, the rollout stage dominates the overall training runtime—for coding tasks, it accounts for over 80%—whereas in some video understanding tasks it accounts for only about 30-40%, a substantial part of this is attributable to the sequence length distribution discussed above. Because the load distribution shifts during training, the time shares of RL stages can be dynamic, especially in the video understanding task. For example, at step 62, there is a sharp spike in rollout time that directly alters the overall stage-wise breakdown. This spike arises from the sudden occurrence of long-tail samples, which induce fluctuations in rollout time. According to Amdahl’s law[1], we should prioritize optimizing the rollout phase. However, the complexity and diversity of the distributions simultaneously pose significant challenges, and also provide opportunities.
Figure 21: Fine-grained time breakdown across different RL tasks. We conduct a detailed profiling of the four tasks on Cluster B, for readability, we focus our in-depth analysis on the rollout and training phases, which typically dominate the overall training time; moreover, the computational characteristics of the inference phase are similar to those of training.
Fine-grained time breakdown. As shown in Figure 21, we perform a finer-grained analysis of the two primary stages—rollout and training. The results reveal substantial cross-task differences in runtime composition, reflecting the underlying sequence length distributions. For the programming task, for instance, decoding accounts for 85.4% of the time, whereas it is only 11.7% for the video understanding task. This disparity directly correlates with the output-length distribution; longer outputs yield a larger decoding share. For tasks with long input sequences (e.g., video understanding), the rollout phase incurs higher communication overhead, primarily due to greater activation exchange of tensor parallel in the prefill phase.
Moreover, we observe considerable scheduling overhead during rollout. The root cause is that contemporary RL frameworks commonly employ inference engines as rollout backend; these engines are designed for online serving scenarios such as iteration-level scheduling and memory management, but these strategies introduce significant CPU overhead. Underscore the need for low CPU overhead inference backends.
In the training stage, we likewise observe substantial communication overhead. For example, in the video understanding task, communication accounts for 32.5% of training time. Two factors primarily contribute: (i) a suboptimal parallelization strategy (here, a simple FSDP setup) that fails to sufficiently hide communication behind computation; and (ii) waiting delays in collective operations due to load imbalance, as illustrated in Figure 20. In the multimodal understanding task, heterogeneity in image and video sizes further exacerbates compute imbalance, indicating that sample-level load-balancing schemes are too coarse-grained.
Tool calling latency is likewise unstable. As illustrated in Figure 22. In a searching task, the searching cost correlates with the difficulty of the query varies a lot. In tool call tasks, some invocations run locally (e.g., calling a Python function, local content retrieval) and are relatively stable, but many rely on remote services (e.g., E2B[6], AI applications[39], model endpoints), whose latency is affected by complex conditions such as network latency, making them inherently more volatile. From the perspective of the RL training framework, asynchronous tool invocation can hide tool latency, and employing caching can mitigate the impact of network instability. Additionally, these remote services entail considerable costs, rendering cost reduction [63] in RL training a compelling area for investigation.
Figure 22: Histogram and CDF search tool calling latency of Searching Task.
Inefficient data transmission and management implementation. Contemporary RL frameworks often adopt a single controller design[60], whereby all inter-stage data transfers must traverse the controller node. This mainly leads to 2 issues: first, in multimodal and long-context tasks, data volumes can be orders of magnitude larger than in standard tasks, making collection and dispatch via the controller prohibitively expensive. For example, in our measurements in cluster A on the current Verl implementation, Ray’s costly serialization/deserialization yields only approximately 0.5 GB/s effective bandwidth within a node and roughly 0.015 GB/s across nodes. Second, funneling all data through a single controller can cause CPU out-of-memory (OOM) on a single node (Sec. 4.5), resulting in training failures. There is a pressing need for a more efficient mechanism to transfer and manage inter-stage data in RL training.
Suboptimal parameter synchronization mechanism. Parameter synchronization further involves cross-framework interaction and becomes increasingly complex and time-consuming in large model training, where there remains room for optimization. Mainstream open-source approaches include HybridEngine, from-disk updates, and all-gather plus broadcast. HybridEngine[60] co-locates rollout and training, simplifying parameter synchronization. The from-disk approach first persists updated parameters, after which the inference engine reloads them from the file system. It is really slow but suitable for auto scaling training. The all-gather plus broadcast scheme first all-gathers parameters to rank 0, which then broadcasts them to the inference engine’s tensor-parallel (TP) group[15]. These approaches either incur redundant communication or lack generality, highlighting the need for a unified, efficient parameter synchronization mechanism.
4.3 GPU Memory Footprint
Figure 23: Memory snapshot[49] of DeepCoder’s training stage. Since each task exhibits similar GPU memory usage during the training stage, we use the programming task for illustration. We offload optimizer states to host memory to prevent OOM.
During RL training, multiple model parameters are involved. For models that are being trained (e.g., the actor model), additional storage must be allocated for gradients and optimizer states[25]. Furthermore, for models that run on the inference engine, sufficient KV-cache storage must also be provisioned[28]. These memory consumers do not coexist throughout the entire process, posing both spatial and temporal challenges for memory management.
Static, coarse-grained memory optimization during training. In RL training, conservative GPU memory policies are often adopted to prevent OOM: static memory footprints (e.g., optimizer states) are preferentially offloaded to host memory, and parameters and gradients are offloaded when device memory becomes insufficient[54, 56]. In addition, as indicated by certain spikes in Figure 23, the long sequence lengths during training lead to substantial activation footprints[32, 20], for which recomputation[19] is commonly employed. Such coarse-grained, static strategies result in suboptimal GPU memory utilization and extra computation or transfer overheads, thereby degrading RL training efficiency.
Load-unaware KV-cache preemption policy. At scales of 16 GPUs, the programming task exhibits frequent KV-cache preemption with tensor parallelism = 8. This arises because the task generally produces long sequences, demanding substantial GPU memory for KV-cache storage. These preemptions incur KV-cache recomputation and directly degrade training efficiency. While existing KV-cache preemption optimizations target inference serving[59, 80], RL training differs substantially and remains largely unexplored.
4.4 Scale Study
Figure 24: Scalability performance across different RL tasks. Experiments were conducted on cluster A. Split refers to dividing the GPU resources evenly into two parts: one part is allocated to the rollout phase, and the other to the inference and training phases.
Figure 25: GPU SM utilization of RL training a 32B LLM using the DAPO algorithm over 32 GPUs.
In this section, we examine the strong scalability of the mainstream RL training framework across different tasks. As shown in Figure 24, we find that different tasks exhibit distinct strong-scaling behavior due to variations in model size and workload distribution. None of the tasks exhibits strong scalability; some tasks (e.g., mathematics) even display negative scaling. To provide an explanation for this observation, we outline four potential factors:
(1) Large GPU idling during rollout. As analyzed in prior sections, the long-tail output phenomenon, where a handful of extremely long sequences appear only on a few inference ranks, forces the remaining ranks to idle until they finish. As depicted in the Figure 25, in the latter portion of the rollout phase, the GPU streaming multiprocessor (SM) utilization for rank 0 reaches 0, whereas rank 24 remains active.
(2) Poor GPU utilization in the rollout phase. During the 2730–3030,s interval shown in Figure 25, the rollout stage is dominated by decoding, which fails to fully exploit GPU compute capacity. Moreover, the decode phase interleaves CPU scheduling, creating short but frequent GPU idles.
(3) High communication cost. As the cluster scales, the aggregate tensor-parallel (TP) communication overhead—especially in the prefill phase—grows and remains uncompensated by overlap. In the searching task rollout stage, 128 GPUs require 585.54,s versus 492.26,s for 64 GPUs (Figure 24). An analogous reversal is observed in training: the math task completes in 301.35,s with 32 GPUs, yet increases to 448.93,s when scaled to 128 GPUs. The reason is that the training stage becomes communication-bound at scale.
4.5 Hyper Parameters Study
Due to the importance of hyperparameters in system optimization[18, 50], in this section, we examine hyper parameters affection to RL training system throughput. Specifically, we discuss the effects of batch size, maximum response length, and, in asynchronous RL training, the maximum permitted staleness on training throughput.
Figure 26: The effect of hyperparameters on different RL tasks. Experiment conducted on cluster B. In Figure (a), execution fails when the batch size is 512 due to a CPU OOM error, while in Figure (b), the failure is caused by the output length exceeding the model’s maximum context length.
Batch size (bsz). In LLM training, batch size is often pivotal for system performance; larger batches can better utilize GPU compute, create opportunities to hide communication overhead, and scale to huge systems[43]. In inference, larger bsz also offers more scheduling opportunities[87, 84]. However, as shown in the Figure 26 (a), the impact of batch size on iteration time in RL training different tasks is affected by bsz in distinct ways, with the following characteristics:
(1) The sensitivity to bsz varies across tasks and is strongly influenced by the model size and the input/output length distributions used by each task. For example, in the searching and video understanding tasks, even with comparable model sizes, differing input/output distribution characteristics lead to markedly different growth trends. In the math task, the extremely long-tailed distribution renders the effect of bsz on iteration time irregular.
(2) At large bsz, CPU out of memory emerges readily in math and video understanding tasks with different reasons. In math, offloading model parameters, gradients, and optimizer states to the CPU leads to excessive memory consumption. In video understanding task, the large bsz induces massive input data (e.g., videos), exhausting memory. This again highlights the importance of storage/memory management mechanisms in RL training.
(3) Notably, in the video understanding task, we observe that training throughput degrades as bsz increases (iteration time increases from 116.86 s at bsz = 128 to 253.40 s (about 2.2×) at bsz = 256). This aligns with our earlier analysis of load imbalance in multimodal tasks. A larger bsz implies more mini-batches, and each mini-batch can suffer imbalance, forcing collective operations to wait for the slowest rank.
Maximum response length. As shown in Figure 26 (b), the impact of the maximum response length on iteration time varies substantially across tasks. The video understanding and searching tasks are almost unaffected by the maximum response length, whereas programming and math exhibit a clear positive correlation. These observations are consistent with our earlier analysis of sequence-length distributions. They also suggest optimization opportunities, such as selecting the maximum response length at the prompt level and adopting task-aware scheduling strategies.
Maximum Permitted Staleness Math500[33] Aime24 Aime23 E2E Time
0 84.4% 32.6% 74.3% 52h 14min
8 84.7% 32.5% 73.0% 33h 50min
16 82.6% 32.7% 74.1% 35h 15min
Table 3: Results under different maximum permitted staleness. Conduct official DeepSeek-R1-Distill-Qwen-1.5B experiment running 80 training steps. For evaluation, we sample 32 responses per question, reporting the average pass@1 accuracy. Experiments are conducted on cluster B, one node for the rollout stage and one for the training stage.
Maximum Permitted Staleness. As shown in Figure 3, asynchronous RL mitigates load imbalance in the rollout phase by placing rollout and training on different machines and executing them in a pipeline. As our analysis shows, rollout time often exceeds training time. Consequently, an asynchronous setup is commonly adopted in which the inference stack uses stale parameters that lag behind the training stage. The degree of staleness can be tuned via the maximum permitted staleness.
As shown in Table 3, asynchronous training delivers roughly a 60% improvement in end-to-end (E2E) throughput over synchronous training, without degrading downstream task performance. Interestingly, the maximum permitted staleness values of 8 and 16 yield similar speedups. This is because, with two nodes, pipeline balance is achieved at around 1–2, so the upper bound is not the limiting factor.
Furthermore, as training scales up, deciding how to allocate GPU nodes between the rollout and training stages to achieve pipeline balance becomes a key systems question. Optimizing this allocation should explicitly account for the maximum permitted staleness in order to identify the optimal configuration. Moreover, the dynamism and heterogeneity of RL workloads also affect resource partitioning and call for adaptive rebalancing. Effective asynchronous RL training requires co-optimizing resource allocation with respect to cluster scale, the maximum permitted staleness, and the distribution of the workload.
5 PolyTrace Benchmark Suite
Motivation: From a research perspective, the benchmark suite serves as a crucial tool for uncovering system insights (Section 4) and enabling researchers to validate their optimization methods effectively. Current RL training systems face challenges in new cases that are often difficult to identify and analyze without comprehensive workload characterization. PolyTrace benchmark suite bridges this gap by providing diverse and realistic workloads from PolyTrace. Also, with access to standardized workloads and performance baselines, researchers can focus on developing novel optimization strategies(e.g., dynamic parallel strategies[72, 12]) rather than spending extensive GPU hours collecting workload data, accelerating the pace of research innovation in RL training systems. From a system evaluation perspective, A benchmark suite provides a standardized platform for comprehensive assessment of RL training frameworks. The benchmark suite enables fair and consistent comparison across different RL training frameworks by providing standardized workloads and evaluation metrics, helping practitioners make decisions about framework selection.
Figure 27: Benchmark Suite Overview.
Benchmark suite design: Our benchmark suite can seamlessly integrate trace data into complex RL training frameworks, consisting of three main processes:
Tool use latency sampler: Based on our previous analysis, the tool invocation time distribution in RL training is highly imbalanced. To simulate real-world scenarios, we first collect tool use times, then perform random sampling according to this latency distribution, enabling the rollout phase to avoid calling actual tools, which often incur significant deployment and invocation costs (e.g., Google’s search API costs \$5 for 1000 calls). This allows us to obtain end-to-end performance of RL tasks in tool invocation scenarios.
Workload sampler and generator: Based on our previous analysis, diverse and real-world workloads are critical for RL training system evaluation as they reveal different performance bottlenecks and ensure comprehensive assessment across realistic training scenarios. To generate workloads, we first sample the specified amount of metadata (including input data length, output data length, and number of multi-turn conversation rounds, etc.) from our trace data based on user-provided hyper parameters (including batch size, task type, step, etc.), then generate dummy inputs according to the data format specified in the RL training framework (such as DataProto in Verl) and put it to the RL training system.
RL system benchmark: After receiving the dummy data generated by the workload sampler and generator, the RL training framework can begin executing the training process. In the rollout phase, we can control the output length of data by setting ignore_eos_token=True and max_output_len=output length for each piece of data, which is widely supported in current inference frameworks, to obtain the actual execution time. Then, the updated dummy data is collected through the inter-stage management backend (such as Ray[41]) and undergoes optimization strategies such as load balancing[60] before being passed to the inference stage, where normal computational processes can proceed, ultimately yielding benchmark performance data.
End-to-end accuracy: To verify the accuracy of the benchmark suite, we use video understanding workloads as an example to conduct performance testing on the Verl framework. As shown in Figure 28, the input data format (DataProto) in the framework primarily consists of two components: batch and no_tensor_batch, requiring only the input/output length information and the shape of video data, which we obtain from PolyTrace through the workload sampler. We then call random methods to populate the DataProto according to the specified format and data types, and directly input it into the Verl framework. The resulting end-to-end latency is 484.48s, while the actual latency is 511.44s, with an error of 5.3%.
6 Discussion
Scope limitations. Despite our best efforts to analyze LLM RL training systems and workloads, our study still has several limitations. Firstly, although our study covers multiple categories of RL tasks, the rapid evolution of RL algorithms means there are emerging task types we did not include. We believe the workloads we selected are representative and that the insights derived from our in-depth analysis are both valuable and broadly applicable. We leave the characterization of these newer workloads to future work. Secondly, our analysis primarily focuses on the RL framework. Other perspectives, such as fault tolerance, are left for future work. Nevertheless, we believe our methodology—characterizing real workloads in an emerging domain and examining how those workloads stress current systems—can be fruitfully applied to other LLM tasks and systems research.
Continuous benchmark enhancement. As RLVR continues to evolve, the benchmarks presented in this paper may be insufficient to fully capture the workloads of RLVF. In response, we will continually expand the diversity of benchmark tasks and incorporate realistic tool-invocation latencies to keep pace with rapidly advancing algorithms. Moreover, we are upgrading our benchmark suite by adding support for additional frameworks and introducing more fine-grained simulations, such as KV cache reuses. Furthermore, we are exploring promising directions, including the development of a simulator that does not require a real execution environment.
7 Conclusion
We present a practical, workload-centric study of RL for LLMs that combines a comprehensive survey with empirical profiling across diverse tasks. Our analysis exposes core challenges and bottlenecks in RL training, highlighting the outsized impact of workload characteristics on framework performance. By profiling three real-world, state-of-the-art RL workloads and augmenting them with open-source algorithmic tasks, we curate a workload dataset that both informs and motivates researchers—and serves as a principled benchmark for evaluating RL framework efficacy.
Appendix A Representative RL tasks
Task Model size Total Step Dataset
Mathematics 32B[68] 191 DAPO[81]
programming 14B[5] 190 DeepCoder[36]
Searching 7B[68] 147 NQ_hotpotqa[24]
Video Understanding 7B[8] 112 RoboVQA[2]
Image Understanding 235B 46 In house
Mathematics 235B 188 In house
Tool Use 235B 59 In house
Table 4: Summary of PolyTrace.
In recent months, RL algorithm researches have emerged, spanning mathematics, programming, tool use, and multimodal content. We select several representative algorithms to conduct detailed system analysis in Section 4.
Mathematics. To enhance LLM reasoning capabilities, industry practice commonly applies reinforcement learning on math-related datasets. DAPO[81] builds on the GRPO[58] algorithm with dynamic sampling policy and provides a high-quality mathematics dataset. We adopt DAPO as the representative algorithm for mathematical tasks.
Programming. A common approach to endow LLMs with strong programming and reasoning ability is to perform RL using coding datasets. DeepCoder[36] curates high-quality open-source data and applies rigorous filtering. After filtering, the dataset contains 24K high-quality coding problems: 7.5K from TACO Verified[30], 16K from PrimeIntellect’s SYNTHETIC-1[38], and 600 from LiveCodeBench[21]. We use this algorithm to represent programming tasks.
Searching. Using external tools enables LLMs to generate more reasonable and reliable responses. To equip LLMs with tool using capacity, many methods have been proposed to improve LLM competence with external tools—such as web search for answering questions, program execution[7] for producing verifiable outputs, and image-processing utilities. Search R1[24] is a highly influential project in searching RL tasks which merges the NQ[27] and HotpotQA[79] training sets to form a unified dataset for RL training. We select it as the representative one.
Video understanding. To enable LLMs to understand video information, numerous algorithms and datasets[62] aim to strengthen LLM performance across multimodal tasks. In particular, the RoboVQA dataset[2] trains LLMs to comprehend video content and issue instructions for a robot’s next actions. We treat it as the representative for video understanding.
Appendix B Related Work
RL training systems. Spurred by rapid algorithmic progress, many RL training systems have appeared in recent months [15, 71]. Verl[60] proposes a single-controller/multi-controller architecture to ensure both usability and efficiency. RealHF[40] leverages careful resource allocation to avoid expensive communication costs. Some systems [29, 91] conduct inter- and intra-stage optimizations to shrink GPU idle time. Furthermore, previous works[77, 90, 14, 48] model multi-stage RL training as a pipeline for acceleration. Beyond this, algorithm–system co-design, such as partial rollout[69, 66], asynchronous RL training[10, 78, 90] has been explored to mitigate the effects of long-tailed sample distributions. Other acceleration techniques, including load-aware scheduling[90] and elastic scaling[91], have also been explored.
Trace analysis. Prior work has analyzed DL workloads across different organizations [22, 16]. Acme[17] from Shanghai AI Lab presents a multifaceted study of cluster jobs predominantly running LLM workloads, covering utilization and fault tolerance. ServeGen[76] provides a detailed analysis of request patterns in Alibaba’s internal LLM services—including text, multimodal, and reasoning workloads—and designs a system for data desensitization. KV Cache in the Wild[73] analyzes KV-cache traces from Alibaba’s LLM services and proposes a more effective management system. Alibaba’s PAI[75] contributes to this discourse by analyzing the challenges encountered within their clusters from both temporal and spatial viewpoints. Concurrently, MegaScale[23] and ByteScale[11] present ByteDance’s experience in training LLMs using a formidable array of over 10,000.